# Short-duration GRB 250221A Afterglow Driven by Two-Component Jets from merger of compact star

### v1

@@ Abstract
GRB 250221A is a short gamma-ray burst (GRB) at redshift $z=0.768$, with a duration of 1.8 s and no extended emission in either Swift/BAT or Konus-Wind bands. A remarkable re-brightening feature in both optical and X-ray bands was observed at $\sim 0.6$ days after the burst trigger, but no supernova or kilonova signature was detected. The burst properties and empirical correlations or distributions (e.g., duration, spectral hardness, location in the Amati correlation, $\varepsilon$-value, $f_{\rm eff}$ parameter, and physical offset) favor a compact binary merger origin. However, a dense circumburst medium with $n\sim 80\rm~cm^{-3}$, obtained by adopting the energy injection into a jet to interpret the late-time re-brightening is inconsistent with the compact binary merger origin. In this paper, we propose a two-component jet model to explain the multiwavelength afterglow observations of GRB 250221A, in which the relativistic narrow jet ($\rm\theta_{c}\sim 3.8^{\circ}$) produces the prompt and the early decay afterglow emission, while the mildly relativistic wide jet ($\rm\theta_{w}\sim 4.4^{\circ}$) dominates at later times, resulting in the observed re-brightening feature. If this is the case, one can obtain a lower medium density with $n\sim 0.72\rm~cm^{-3}$ which is a little bit higher than that of short GRBs in merger environments, but falls into the reasonable and acceptable range. Finally, a possible kilonova emission is also discussed within the scenario of compact star merger origin of GRB 250221A.
@@ I Introduction
Gamma-ray bursts (GRBs) are among the most energetic transients in the Universe. Observationally, based on the duration of prompt gamma-ray emission (e.g., $T_{\rm 90}$), GRBs are divided into two categories: long-duration GRBs (LGRBs) with $T_{\rm 90}>2$ s, and short-duration GRBs (SGRBs) with $T_{\rm 90}<2$ s (Kouveliotou et al., 1993; Bromberg et al., 2013). Several LGRBs have been confirmed to associate with broad-lined type Ic supernovae (SNe), and the progenitor of LGRBs are generally believed to be the collapse of massive stars (Galama et al., 1998; Malesani et al., 2004; Modjaz et al., 2006; Pian et al., 2006; Zhang, 2006; Kelly & Kirshner, 2012; Lü et al., 2018). In contrast, mergers of compact binary star objects, such as neutron star-neutron star (NS–NS), or neutron star–black hole (NS–BH), are considered as possible progenitors of SGRBs (Zhang, 2006; Jin et al., 2016; Lü et al., 2017; Lamb et al., 2019; Metzger, 2019). The “smoking gun” evidence is SGRB 170817A, in which both the kilonova (KN) electromagnetic emission AT2017gfo, powered by radioactive decay from r-process nuclei, and the corresponding gravitational-wave (GW) source GW170817 were simultaneously detected (Abbott et al., 2017; Metzger, 2019). Whether it is the collapse of a massive star or the merger of compact stars, a hyperaccreting black hole (Popham et al., 1999; Lei et al., 2013; Liu et al., 2017), or a rapidly spinning, strongly magnetized NS, named a millisecond magnetar (Usov, 1992; Dai & Lu, 1998a, b; Zhang & Mészáros, 2001; Metzger et al., 2011; Bucciantini et al., 2012; Lü & Zhang, 2014) can be formed as the central engine of GRBs.
@@
The central engine drives an ultrarelativistic jet, which releases energy through internal shocks or magnetic reconnection to produce prompt $\gamma$-ray emission (Meszaros & Rees, 1993; Rees & Meszaros, 1994), while the interaction of the jet with the external medium generates multiwavelength afterglow emission (Mészáros & Rees, 1997; Sari et al., 1998). The structure of ultrarelativistic jets was traditionally considered as a top-hat, in which the energy density and Lorentz factor are constant within a core opening angle (Rhoads, 1997; Sari et al., 1999). However, results of numerical simulations and observations of SGRB 170817A reveal that jets often possess an angular structure, such as a Gaussian distribution (Zhang & Mészáros, 2002a; Kumar & Granot, 2003; Zhang et al., 2004), a power-law distribution (Mészáros et al., 1998; Rossi et al., 2002), or two-component jets (Huang et al., 2004; Peng et al., 2005; Racusin et al., 2008; Filgas et al., 2011). In addition, when an ultrarelativistic jet crosses the stellar envelope or the merger ejecta, it can form a jet–cocoon system (Nakar & Piran, 2017; Lazzati & Perna, 2019; Gottlieb et al., 2018), whose low-energy, wide-angle emission is particularly prominent in off-axis observations.
@@
From the observational point of view, the behavior of re-brightening in X-ray and optical afterglows has been detected with peak times typically occurring between $10^{\rm 2}-10^{\rm 5}$ seconds (Liang et al., 2013; Yang et al., 2024). The physical interpretation of such a re-brightening feature is to invoke a structured jet with off-axis (Nakar et al., 2003; Berger et al., 2003; Huang et al., 2004), or late energy injection from slower ejecta or central engine activity (refreshed shock) (Zhang & Mészáros, 2002b; Björnsson et al., 2004; Laskar et al., 2015), or a medium density jump (Dai & Lu, 2002; Lazzati et al., 2002; Dai & Wu, 2003), or the existence of two component jets (Huang et al., 2004; Wu et al., 2005; Racusin et al., 2008).
@@
Recently, a short-duration GRB 250221A, with a measured redshift of $z=0.768$ was detected, and the late-time afterglow of this GRB exhibited an unusual re-brightening feature approximately 0.6 days post-burst. The burst properties (duration, spectral hardness, peak energy, and location in the Amati relation) all favor a compact binary merger origin (Angulo-Valdez et al., 2026), and suggest an internal energy injection from a long-lasting refreshed shock into the jet. However, it is found that the medium density of the circumburst is as high as $n\sim 80\rm~cm^{-3}$, which is inconsistent with the typical density of the environment of merger systems (Zhang, 2018). So, Angulo-Valdez et al. (2026) claimed that the GRB 250221A is likely originated from the scenario of massive star collapse. In this paper, by comparing the properties of GRB 250221A (e.g., $\varepsilon$ value, $f_{\rm eff}$ parameter, physical offset, and star formation rate) with other typical LGRBs and SGRBs, we propose a two-component jet model to explain the multiwavelength afterglow observations of GRB 250221A. It is found that GRB 250221A is consistent with the scenario of compact star mergers.
@@
The paper is organized as follows. The observations of the prompt and afterglow emission of the short GRB 250221A are presented in Section 2. In Section 3, we compare the properties of GRB 250221A with other long and short GRBs. The physical interpretation with a two-component jet model and the results are shown in Section 4. In Section 5, we attempt to calculate the possible kilonova emission to compare with observations. The conclusions are drawn in Section 6 with some additional discussion. Throughout this paper, we adopt a standard concordance cosmology with $\Omega_{\rm M}=0.31$, $\Omega_{\rm\Lambda}=0.69$, and $H_{0}=67.7~\rm km~\rm s^{-1}~Mpc^{-1}$.
@@ II The Observations and Data Analysis
At 03:34:37 UT on February 21, 2025 (as $T_{0}$), the Burst Alert telescope (BAT) of Swift triggered and located GRB 250221A, which was simultaneously detected by Konus-Wind (Palmer et al., 2025; Frederiks et al., 2025). The light curve of the prompt emission of this burst exhibits a single short-pulse structure with a duration of $T_{\rm 90}=1.8\pm 0.3$ s in 15-350 keV (Palmer et al., 2025), and the prompt $\gamma$-ray spectrum is well fitted by a power-law model with a photon index of $\Gamma=1.43\pm 0.23$ due to its narrow energy band. However, over the wide energy band range of Konus-Wind, the time-integrated spectrum of the burst can be fitted by a power-law with exponential cutoff (CPL) model, with a low-energy index of $\alpha=-1.05\pm 0.12$ and a peak energy of $E_{\rm p}=242^{+26}_{-22}$ keV (Frederiks et al., 2025), and the fluence in the 10 keV -10 MeV band is $(1.67\pm 0.09)\times 10^{-6}\rm~erg~cm^{-2}$.
@@
Several optical telescopes reported the follow-up observations of GRB 250221A, such as the Very Large Telescope (VLT; Palmerio et al. 2025), the DDRAGO wide-field imager on the COLIBRI telescope (Watson et al., 2025), as well as radio data from the Australia Telescope Compact Array (ATCA; Gulati et al. 2025) and the Very Large Array (VLA; Ricci & Troja 2025). We collected the optical data from Angulo-Valdez et al. (2026), and the optical light curve of GRB 250221A shows an initial decay followed by a re-brightening at around 0.6 days. Moreover, Angulo-Valdez et al. (2026) analyzed the early afterglow spectrum of GRB 250221A, in which multiple emission and absorption lines were detected. They reported a redshift of $z=0.768$ for the host galaxy and confirmed that the GRB is located within this galaxy, with a projected offset of $7^{+12}_{-7}$ kpc. They also derived a star formation rate (SFR) of approximately $3~M_{\odot}~\rm yr^{-1}$ for the galaxy.
III Comparison of the Properties of GRB 250221A with Other Long and Short GRBs
Even the duration of a GRB can provide us with information on the timescale of central engine activity, but the duration is sensitively dependent on the observed energy and the instrument (Qin et al., 2013). It suggests that the duration is no longer a reliable indicator of the physical origin of a GRB (Zhang et al., 2009; Lü et al., 2010). For example, several long-duration GRBs, i.e., GRB 060614 (Gehrels et al., 2006; Yang et al., 2015), GRB 211211A (Rastinejad et al., 2022; Troja et al., 2022; Yang et al., 2022; Chang et al., 2023; Gompertz et al., 2023; Liu et al., 2025), GRB 211227A (Lü et al., 2022; Ferro et al., 2023), and GRB 230307A (Dichiara et al., 2023; Levan et al., 2024; Yang et al., 2024; Du et al., 2024), are thought to be from the merger of two compact stars. On the contrary, the short-duration GRB 200826A is believed to originate from the collapse of massive stars (Ahumada et al., 2021; Zhang et al., 2021; Rossi et al., 2022). So, Zhang (2006) proposed adopting Type I (e.g., compact star mergers) and Type II (e.g., death of massive stars) to classify GRBs. If possible, the presence of supernova or kilonova emission associated with GRBs can directly provide evidence to determine the nature of the GRB progenitor (Zhang, 2018). However, catching such associated events remains challenging, especially the kilonova emission. From the statistical point of view, different progenitors of GRBs may follow different empirical correlations or distributions (e.g., Amati relation, $\varepsilon$-value, $f_{\rm eff}$ parameter, physical offset, and star formation rate). In this Section, we compare several properties of GRB 250221A with those of other long and short GRBs.
@@ III.1 $E_{\rm p}-E_{\rm\gamma,iso}$ correlation
Statistically, Amati et al. (2002) discovered a relation between the peak energy (e.g., $E_{\rm p}$) and the isotropic-equivalent energy more energetic (e.g., $E_{\rm\gamma,iso}$) for most long-duration GRBs, and it can be expressed as $E_{\rm p}(1+z)\propto E^{1/2}_{\rm\gamma,iso}$ (called Amati relation). However, most Type I GRBs are inconsistent with Type II GRBs, and it seems to be a little shallower for the power index compared with that of Type II GRBs (Zhang et al., 2009). Therefore, the Amati relation may play an important role in distinguishing between Type II and Type I GRBs. By adopting the values of $E_{\rm p}=242^{+26}_{-22}$ and $E_{\rm\gamma,iso}=(2.8\pm 0.2)\times 10^{51}\rm~erg$ from Konus-Wind observations (Frederiks et al., 2025), we plot GRB 250221A on the Amati relation diagram. It is found that GRB 250221A lies outside the $95\%$ confidence region of Type II GRBs, but seems to be closer to the Type I GRB population (see Figure 1(a)). This indicates that GRB 250221A is more likely to originate from a compact object merger (e.g., NS–NS or NS–BH) rather than from the collapse of a massive star.
III.2 $\varepsilon$-value
@@
where $E_{\rm\gamma,iso,52}$ and $E_{\rm p,z,2}$ are expressed in units of $10^{52}\rm~erg$ and $10^{2}\rm~keV$, respectively. Statistical analysis of GRB samples shows that $\varepsilon$ exhibits a bimodal distribution with $\varepsilon=0.03$ as the dividing line between the high- and low-$\varepsilon$ regions. In this classification method, GRBs located in the high-$\varepsilon$ (e.g., Type II) and low-$\varepsilon$ (e.g., Type I) regions are likely to be associated with the collapse of massive stars and compact star mergers, respectively. By adopting the method, one can obtain a value of $\varepsilon\sim 0.024$ for GRB 250221A, as shown in Figure 1(b). It is found that GRB 250221A is more likely to be located in the low-$\varepsilon$ region, but it is very close to the boundary between the long and short GRB populations. Therefore, GRB 250221A seems to be compatible with a compact object merger origin, but its ambiguity $\varepsilon$-value can not provide strongly evidence to support such origin from merger of compact objects.
III.3 “Tip of iceberg” effect
The duration of GRBs may not be intrinsic to reflect the activity of the GRB central engine. For example, the majority of a GRB emission may fall below the background noise, and only the brightest pulses can be detected. As a result, the event can be observationally misclassified as short GRBs, a phenomenon known as the “tip of iceberg” effect. In order to test this effect, Lü et al. (2014) introduced an amplitude parameter as a third dimension. The amplitude parameter is defined as $f_{\rm eff}=F^{{}^{\prime}}_{\rm p}/F_{\rm B}$, where $F^{{}^{\prime}}_{\rm p}$ denotes the peak flux of a pseudo-GRB, representing an intrinsically identical GRB with a lower amplitude, and $F_{\rm B}$ is the background flux. They found that the values of $f_{\rm eff}$ for most short GRBs are statistically larger than those for long GRBs, and suggested that most short GRBs are intrinsically short without being the “tip of iceberg” of long GRBs (Lü et al., 2014). By extracting the prompt emission light curve of GRB 250221A from Swift/BAT, one can calculate the value of $f_{\rm eff}\sim 1.68$ for GRB 250221A to adopt the same method from Lü et al. (2014). We find that the location of GRB 250221A is consistent with the typical distribution of short GRBs (see Figure 1(c)), and suggest that GRB 250221A is an intrinsically short GRB, rather than due to the “tip of iceberg” effect of long GRBs.

### v2

# arXiv:2602.21516v2[astro-ph.HE] 28 Feb 2026

@@ Abstract
GRB 250221A is a short gamma-ray burst (GRB) at redshift $z=0.768$, with a duration of 1.8 s and no extended emission in either Swift/Burst Alert Telescope or Konus-Wind bands. A remarkable rebrightening feature in both optical and X-ray bands was observed at $\sim$0.6 days after the burst trigger, but no supernova or kilonova signature was detected. The burst properties and empirical correlations or distributions (e.g., duration, spectral hardness, location in the Amati correlation, $\varepsilon$-value, $f_{\rm eff}$ parameter, and physical offset) favor a compact binary merger origin. However, a dense circumburst medium with $n\sim 80\rm~cm^{-3}$, obtained by adopting the energy injection into a jet to interpret the late-time rebrightening is inconsistent with the compact binary merger origin. In this paper, we propose a two-component jet model to explain the multiwavelength afterglow observations of GRB 250221A, in which the relativistic narrow jet ($\rm\theta_{c}\sim 3.8^{\circ}$) produces the prompt and the early decay afterglow emission, while the mildly relativistic wide jet ($\rm\theta_{w}\sim 4.4^{\circ}$) dominates at later times, resulting in the observed rebrightening feature. If this is the case, one can obtain a lower medium density with $n\sim 0.72\rm~cm^{-3}$ which is a little bit higher than that of short GRBs in merger environments, but falls into the reasonable and acceptable range. Finally, a possible kilonova emission is also discussed within the scenario of compact star merger origin of GRB 250221A.
@@ I Introduction
Gamma-ray bursts (GRBs) are among the most energetic transients in the Universe. Observationally, based on the duration of prompt gamma-ray emission (e.g., $T_{\rm 90}$), GRBs are divided into two categories: long-duration GRBs (LGRBs) with $T_{\rm 90}>2$ s, and short-duration GRBs (SGRBs) with $T_{\rm 90}<2$ s (Kouveliotou et al., 1993; Bromberg et al., 2013). Several LGRBs have been confirmed to be associated with broad-lined Type Ic supernovae (SNe), and the progenitor of LGRBs are generally believed to be the collapse of massive stars (Galama et al., 1998; Malesani et al., 2004; Modjaz et al., 2006; Pian et al., 2006; Zhang, 2006; Kelly & Kirshner, 2012; Lü et al., 2018). In contrast, mergers of compact binary star objects, such as neutron star-neutron star (NS–NS), or neutron star–black hole (NS–BH), are considered as possible progenitors of SGRBs (Zhang, 2006; Jin et al., 2016; Lü et al., 2017; Lamb et al., 2019; Metzger, 2019). The “smoking gun” evidence is SGRB 170817A, in which both the kilonova (KN) electromagnetic emission AT2017gfo, powered by radioactive decay from r-process nuclei, and the corresponding gravitational-wave (GW) source GW170817 were simultaneously detected (Abbott et al., 2017; Metzger, 2019). Whether it is the collapse of a massive star or the merger of compact stars, a hyperaccreting BH (Popham et al., 1999; Lei et al., 2013; Liu et al., 2017), or a rapidly spinning, strongly magnetized NS, named a millisecond magnetar (Usov, 1992; Dai & Lu, 1998a, b; Zhang & Mészáros, 2001; Metzger et al., 2011; Bucciantini et al., 2012; Lü & Zhang, 2014) can be formed as the central engine of GRBs.
@@
The central engine drives an ultrarelativistic jet, which releases energy through internal shocks or magnetic reconnection to produce prompt gamma-ray emission (Meszaros & Rees, 1993; Rees & Meszaros, 1994), while the interaction of the jet with the external medium generates multiwavelength afterglow emission (Mészáros & Rees, 1997; Sari et al., 1998). The structure of ultrarelativistic jets was traditionally considered as a top-hat, in which the energy density and Lorentz factor are constant within a core opening angle (Rhoads, 1997; Sari et al., 1999). However, results of numerical simulations and observations of SGRB 170817A reveal that jets often possess an angular structure, such as a Gaussian distribution (Zhang & Mészáros, 2002a; Kumar & Granot, 2003; Zhang et al., 2004), a power-law distribution (Mészáros et al., 1998; Rossi et al., 2002), or two-component jets (Huang et al., 2004; Peng et al., 2005; Racusin et al., 2008; Filgas et al., 2011). In addition, when an ultrarelativistic jet crosses the stellar envelope or the merger ejecta, it can form a jet–cocoon system (Nakar & Piran, 2017; Lazzati & Perna, 2019; Gottlieb et al., 2018), whose low-energy, wide-angle emission is particularly prominent in off-axis observations.
@@
From the observational point of view, the behavior of rebrightening in X-ray and optical afterglows has been detected with peak times typically occurring between $10^{\rm 2}$ and $10^{\rm 5}$ seconds (Liang et al., 2013; Yang et al., 2024). The physical interpretation of such a rebrightening feature is to invoke a structured jet with off-axis (Nakar et al., 2003; Berger et al., 2003; Huang et al., 2004), or late energy injection from slower ejecta or central engine activity (refreshed shock) (Zhang & Mészáros, 2002b; Björnsson et al., 2004; Laskar et al., 2015), or a medium density jump (Dai & Lu, 2002; Lazzati et al., 2002; Dai & Wu, 2003), or the existence of two-component jets (Huang et al., 2004; Wu et al., 2005; Racusin et al., 2008).
@@
Recently, a short GRB 250221A, with a measured redshift of $z=0.768$ was detected, and the late-time afterglow of this GRB exhibited an unusual rebrightening feature $\sim$0.6 days postburst. The burst properties (duration, spectral hardness, peak energy, and location in the Amati relation) all favor a compact binary merger origin (Angulo-Valdez et al., 2026), and suggest an internal energy injection from a long-lasting refreshed shock into the jet. However, it is found that the medium density of the circumburst is as high as $n\sim 80\rm~cm^{-3}$, which is inconsistent with the typical density of the environment of merger systems (Zhang, 2018). So, Angulo-Valdez et al. (2026) claimed that the GRB 250221A is likely originated from the scenario of massive star collapse. In this paper, by comparing the properties of GRB 250221A (e.g., $\varepsilon$ value, $f_{\rm eff}$ parameter, physical offset, and star formation rate (SFR)) with those of other typical LGRBs and SGRBs, we propose a two-component jet model to explain the multiwavelength afterglow observations of GRB 250221A. It is found that GRB 250221A is consistent with the scenario of compact star mergers.
@@
The paper is organized as follows. The observations of the prompt and afterglow emission of the SGRB 250221A are presented in Section 2. In Section 3, we compare the properties of GRB 250221A with those of other long and short GRBs. The physical interpretation with a two-component jet model and the results are shown in Section 4. In Section 5, we attempt to calculate the possible kilonova emission to compare with observations. The conclusions are drawn in Section 6 with some additional discussion. Throughout this paper, we adopt a standard concordance cosmology with $\Omega_{\rm M}=0.31$, $\Omega_{\rm\Lambda}=0.69$, and $H_{0}=67.7~\rm km~\rm s^{-1}~Mpc^{-1}$.
@@ II The Observations and Data Analysis
At 03:34:37 UT on 2025 February 21 (as $T_{0}$), the Burst Alert Telescope (BAT) of Swift triggered and located GRB 250221A, which was simultaneously detected by Konus-Wind (Palmer et al., 2025; Frederiks et al., 2025). The light curve of the prompt emission of this burst exhibits a single short-pulse structure with a duration of $T_{\rm 90}=1.8\pm 0.3$ s in 15-350 keV (Palmer et al., 2025), and the prompt gamma-ray spectrum is well fitted by a power-law model with a photon index of $\Gamma=1.43\pm 0.23$ due to its narrow energy band. However, over the wide energy band range of Konus-Wind, the time-integrated spectrum of the burst can be fitted by a power-law with exponential cutoff (CPL) model, with a low-energy index of $\alpha=-1.05\pm 0.12$ and a peak energy of $E_{\rm p}=242^{+26}_{-22}$ keV (Frederiks et al., 2025), and the fluence in the 10 keV -10 MeV band is $(1.67\pm 0.09)\times 10^{-6}\rm~erg~cm^{-2}$.
@@
Several optical telescopes reported the follow-up observations of GRB 250221A, such as the Very Large Telescope (VLT; Palmerio et al. 2025), the DDRAGO wide-field imager on the COLIBRI telescope (Watson et al., 2025), as well as radio data from the Australia Telescope Compact Array (ATCA; Gulati et al. 2025) and the Very Large Array (VLA; Ricci & Troja 2025). We collected the optical data from Angulo-Valdez et al. (2026), and the optical light curve of GRB 250221A shows an initial decay followed by a rebrightening at around 0.6 days. Moreover, Angulo-Valdez et al. (2026) analyzed the early afterglow spectrum of GRB 250221A, in which multiple emission and absorption lines were detected. They reported a redshift of $z=0.768$ for the host galaxy and confirmed that the GRB is located within this galaxy, with a projected offset of $7^{+12}_{-7}$ kpc. They also derived an SFR of approximately $3~M_{\odot}~\rm yr^{-1}$ for the galaxy.
III Comparison of the Properties of GRB 250221A with Those of Other Long and Short GRBs
Even the duration of a GRB can provide us with information on the timescale of central engine activity, but the duration is sensitively dependent on the observed energy and the instrument (Qin et al., 2013). It suggests that the duration is no longer a reliable indicator of the physical origin of a GRB (Zhang et al., 2009; Lü et al., 2010). For example, several LGRBs, i.e., GRB 060614 (Gehrels et al., 2006; Yang et al., 2015), GRB 211211A (Rastinejad et al., 2022; Troja et al., 2022; Yang et al., 2022; Chang et al., 2023; Gompertz et al., 2023; Liu et al., 2025), GRB 211227A (Lü et al., 2022; Ferro et al., 2023), and GRB 230307A (Dichiara et al., 2023; Levan et al., 2024; Yang et al., 2024; Du et al., 2024), are thought to be from the merger of two compact stars. On the contrary, the SGRB 200826A is believed to originate from the collapse of massive stars (Ahumada et al., 2021; Zhang et al., 2021; Rossi et al., 2022). So, Zhang (2006) proposed adopting Type I (e.g., compact star mergers) and Type II (e.g., death of massive stars) to classify GRBs. If possible, the presence of supernova or kilonova emission associated with GRBs can directly provide evidence to determine the nature of the GRB progenitor (Zhang, 2018). However, catching such associated events remains challenging, especially the kilonova emission. From the statistical point of view, different progenitors of GRBs may follow different empirical correlations or distributions (e.g., Amati relation, $\varepsilon$ value, $f_{\rm eff}$ parameter, physical offset, and SFR). In this section, we compare several properties of GRB 250221A with those of other long and short GRBs.
@@ III.1 $E_{\rm p}-E_{\rm\gamma,iso}$ correlation
Statistically, Amati et al. (2002) discovered a relation between the peak energy (e.g., $E_{\rm p}$) and the isotropic-equivalent energy more energetic (e.g., $E_{\rm\gamma,iso}$) for most LGRBs, and it can be expressed as $E_{\rm p}(1+z)\propto E^{1/2}_{\rm\gamma,iso}$ (called the Amati relation). However, most Type I GRBs are inconsistent with Type II GRBs, and it seems to be a little shallower for the power index compared with that of Type II GRBs (Zhang et al., 2009). Therefore, the Amati relation may play an important role in distinguishing between Type II and Type I GRBs. By adopting the values of $E_{\rm p}=242^{+26}_{-22}$ and $E_{\rm\gamma,iso}=(2.8\pm 0.2)\times 10^{51}\rm~erg$ from Konus-Wind observations (Frederiks et al., 2025), we plot GRB 250221A on the Amati relation diagram. It is found that GRB 250221A lies outside the $95\%$ confidence region of Type II GRBs, but seems to be closer to the Type I GRB population (see Figure 1(a)). This indicates that GRB 250221A is more likely to originate from a compact object merger (e.g., NS–NS or NS–BH) rather than from the collapse of a massive star.
III.2 $\varepsilon$ value
@@
where $E_{\rm\gamma,iso,52}$ and $E_{\rm p,z,2}$ are expressed in units of $10^{52}\rm~erg$ and $10^{2}\rm~keV$, respectively. Statistical analysis of GRB samples shows that $\varepsilon$ exhibits a bimodal distribution with $\varepsilon=0.03$ as the dividing line between the high- and low-$\varepsilon$ regions. In this classification method, GRBs located in the high-$\varepsilon$ (e.g., Type II) and low-$\varepsilon$ (e.g., Type I) regions are likely to be associated with the collapse of massive stars and compact star mergers, respectively. By adopting the method, one can obtain a value of $\varepsilon\sim 0.024$ for GRB 250221A, as shown in Figure 1(b). It is found that GRB 250221A is more likely to be located in the low-$\varepsilon$ region, but it is very close to the boundary between the long and short GRB populations. Therefore, GRB 250221A seems to be compatible with a compact object merger origin, but its ambiguous $\varepsilon$ value cannot provide strong evidence to support such an origin from the merger of compact objects.
III.3 “Tip of the iceberg” effect
The duration of GRBs may not be intrinsic to reflect the activity of the GRB central engine. For example, the majority of a GRB emission may fall below the background noise, and only the brightest pulses can be detected. As a result, the event can be observationally misclassified as short GRBs, a phenomenon known as the “tip of the iceberg” effect. In order to test this effect, Lü et al. (2014) introduced an amplitude parameter as a third dimension. The amplitude parameter is defined as $f_{\rm eff}=F^{{}^{\prime}}_{\rm p}/F_{\rm B}$, where $F^{{}^{\prime}}_{\rm p}$ denotes the peak flux of a pseudo-GRB, representing an intrinsically identical GRB with a lower amplitude, and $F_{\rm B}$ is the background flux. They found that the values of $f_{\rm eff}$ for most short GRBs are statistically larger than those for long GRBs, and suggested that most short GRBs are intrinsically short without being the “tip of iceberg” of long GRBs (Lü et al., 2014). By extracting the prompt emission light curve of GRB 250221A from Swift/BAT, one can calculate the value of $f_{\rm eff}\sim 1.68$ for GRB 250221A to adopt the same method from Lü et al. (2014). We find that the location of GRB 250221A is consistent with the typical distribution of short GRBs (see Figure 1(c)), and suggest that GRB 250221A is an intrinsically short GRB, rather than due to the “tip of the iceberg” effect of long GRBs.